# Cartesian atomic cluster expansion for machine learning interatomic potentials

### v2

@@ III.1 Bulk water
For running molecular dynamics (MD) simulations using MLIPs, besides accuracy, a crucial requirement is the stability - the possibility to run long-enough trajectories with the system remaining in physically reasonable states and the simulation not exploding, even at conditions with few or no training data. To assess the simulation stability, we performed independent simulations of 300 ps in length for water at 1 g/mL at 300 K, 500 K, 1000 K, 1500 K, and 2000 K using the CACE $T=1$ model. The simulation cell contains 512 water molecules. The time step was 1 fs, and Nosé–Hoover thermostat was used. The upper panel of Fig. 3 shows the oxygen-oxygen (O-O) radial distribution function (RDF). At 300 K, the computed O-O RDF is in excellent agreement with experiment from X-ray diffraction measurements Skinner et al. (2014). Note that the nuclear quantum effects have a slight destructuring effect in the O-O RDF, but the effect is quite small Cheng et al. (2019). The mean squared displacements (MSD) from these simulations are shown on the lower panel of Fig. 3. The diffusivity $D$, obtain by fitting a linear function to the MSD, at 300 K agrees well with the DFT MD result ($D_{\mathrm{DFT}}=2.67\pm 0.10\mathrm{\AA}^{2}/$ps) from Ref. Marsalek and Markland (2017). At high temperatures up to 2000 K, the MD simulations remained stable, which demonstrate the superior stability of the CACE potential. Note that we do not expect the result at the very high temperatures to be physically predictive, and this example is just to illustrate the stability of CACE. Such stability is not only necessary for converging statistical properties, but also allows the refinement of the potential: to iteratively improve the accuracy of the potential under certain conditions, one can perform active learning which collects snapshots from corresponding MD trajectories and use these snapshots for re-training.
III.2 Small molecules: ethanol and 3BPA
To demonstrate CACE on small organic molecules, we fitted to the ethanol sampled from DFT molecular dynamics simulations at 500 K, which is from the MD17 database Chmiela et al. (2017) that is often used for benchmark purposes. We used a cutoff $r_{\mathrm{cut}}=4$, 6 Bessel radial functions, $c=12$, $l_{\mathrm{max}}=4$, $\nu_{\mathrm{max}}=3$, $N_{\mathrm{embedding}}=4$, and one message passing layer ($T=1$). The model was trained on 1,000 structures randomly selected from the dataset (90% train/10% validation split), and tested on another 10,000 random structures.
Refer to captionaRefer to caption Refer to captionb
Figure 4: The dihedral scan benchmark of the 3BPA molecule. a Molecular structure and three dihedral angles. b The dihedral potential energy landscape for $\beta=120^{\circ}$, as predicted by DFT and CACE. The white dots on the DFT potential energy surface correspond to all the training configurations that have $\beta$ between ${}^{\circ}$ and ${}^{\circ}$.
@@ III.2 Small molecules: ethanol and 3BPA
Table 2: Mean absolute error (MAE) of energy (E) and force (F) for ethanol in the MD-17 dataset, in units of meV and meV/Å, respectively, trained on 1000 reference configurations. The stability (S) in the unit of ps shows how long the simulation using the MLIP is physically realistic in MD of a maximum length of 300 ps at 500 K. Besides CACE, the errors are obtained from respective references, and the stability is from Ref. Fu et al. (2023).
Table 2 shows a comparison of the different MLIPs trained on energies and forces for the same MD17-ethanol dataset. The energy accuracy of CACE is similar to the NequIP model with $T=5$, although the CACE force error is larger. The stability (S) in Table 2 measures how long MD simulations at 500 K using the MLIP is physically sensical (no pathological behavior or enter physically prohibitive states) in runs with a maximum length of 300 ps, so 300 ps is the maximum score in this case Fu et al. (2023). The stability values for other MLIPs are from Ref. Fu et al. (2023) which used 10,000 training structures and a time step of 0.5 fs in MD. For CACE, we used 1,000 training structures and a time step of 1 fs, which is more stringent, but the MD had nevertheless remained stable. As observed in Refs. Fu et al. (2023); Stocker et al. (2022), force accuracy does not imply stability, and the latter is a key metric for MLIPs.
@@
Table 3: Root mean square errors (RMSE) for energy (E) and force (F) on the 3BPA temperature transferability dataset, reported in units of meV and meV/Å. All models were trained on $T=300$ K data. Results of previous models except for NequIP and MACE are from Kovács et al. (2021). ANI-2x was trained on 8.9 million structures, and ANI-pretrained was pretrained and fine-tuned on this dataset.
As another example, we fitted to the 3BPA (3-(benzyloxy)pyridin-2-amine) dataset Kovács et al. (2021) consisting of 500 training structures at 300 K. Test data are from simulations at 300 K, 600 K, and 1200 K. We used $r_{\mathrm{cut}}=6$, 8 Bessel radial functions, $c=12$, $l_{\mathrm{max}}=4$, $\nu_{\mathrm{max}}=3$, $N_{\mathrm{embedding}}=4$, and $T=1$. Table 3 shows the errors of different potentials when extrapolating to higher temperature. NequIP and MACE perform the best, while CACE is somewhere between them and ACE. We then compared the dihedral potential energy surface of the molecule. Fig. 4 shows the complex energy landscapes of the $\alpha$-$\gamma$ plane predicted by DFT and by CACE, for the case of $\beta=120^{\circ}$, the plane with the fewest training points. Nevertheless, the CACE landscape closely resembles the ground truth and is regular, more so than the other potentials benchmarked in Ref. Kovács et al. (2021) despite those were trained on a more diverse dataset including high temperature configurations.
III.3 25-element high-entropy alloys
High-entropy alloys, which are composed of several metallic elements, have unique mechanical, thermodynamics, and catalytic properties George et al. (2019). They are challenging to model due to the high number of elements involved, for which many MLIPs scale poorly with. Ref. Lopanitsyna et al. (2023) introduced 25 d-block transition metal HEA dataset that contains 25,630 distorted crystalline structures containing 36 or 48 on bcc or fcc lattices. Ref. Lopanitsyna et al. (2023) further performed an alchemical learning (AL) that explicit compresses chemical space, leading to a model that is both interpretable and very stable.
Figure 5: Benchmark on the HEA25 dataset. a Learning curves for different models. Both alchemical learning (AL) models are from Ref. Lopanitsyna et al. (2023). b The first two principal components (PCs) of the CACE embedding matrix. The periods are highlighted with orange, blue, and green lines. Interpolated positions for Re and Os are indicated with empty circles. c Energy comparison between DFT and CACE for a high temperature (5000 K) and a low temperature (300 K) trajectory. The inset shows example solid and melt configurations from the 300 K and the 5000 K trajectories, respectively. d Comparison between the substitution energy $E_{\mathrm{sub}}$, the potential energy difference between the original and the structure substituted with Re and Os element per substitution atom, computed using DFT and CACE. The inset shows the parity plot for the force components computed for those structures.
We took a test set containing 1,000 configurations, and used up to the rest of the 24,630 data points for training and validation. We used $r_{\mathrm{cut}}=4.5$, $l_{\mathrm{max}}=3$, $\nu_{\mathrm{max}}=3$, $T=1$, and $N_{\mathrm{embedding}}=4$ for aggressive alchemical compression.
@@
Table 4: Mean absolute error (MAE) of energy (E) and force (F) for the HEA25 dataset, reported in units of meV and meV/Å. There are three sets of benchmarks on test error, alchemical extrapolation to unseen Re and Os substitute atoms, and temperature transferability to 5000 K.
@@
A stringent way to demonstrate alchemical learning is the extrapolation to unseen elements. In this case, Re and Os are absent from the training set. Ref. Lopanitsyna et al. (2023) provides a hold-out set containing 61 pairs of structures, each of 36 or 48 atoms. Each pair has one fcc or bcc structure with only the original 25 elements, and another structure with up to 11 atoms randomly substituted with Re and Os. We simply took the CACE model trained for the 25-element dataset, and obtained the chemical embedding of Re and Os by a linear interpolation between W and Ir, as illustrated by the empty circles in Fig. 5b. We also added atomic-energy baselines for Re and Os (obtained by fitting the residual energy to a two-parameter linear model that only depends on the Re and Os content). In Fig. 5d we show the parity plot for the substitution energy, defined as the potential energy difference between the original and the substituted structure per substitution atom. Impressively, the CACE error for those substituted structures with unseen elements are rather similar with the test error on the 25-element training set, and is lower than the AL method in Ref. Lopanitsyna et al. (2023). The force comparison, shown in the inset of Fig. 5d, shows the same trend. With the PET potential, it is not clear whether alchemical learning can be performed in a straightforward way.
@@ IV Discussion and limitations
Atomic cluster expansion and equivariant MPNNs are probably both amongst the biggest breakthroughs in the field of MLIPs: the ACE formulation not only guarantees a complete basis for describing atomic environments, but also provides a framework for understanding and connecting different descriptors Drautz (2019); Dusson et al. (2022). E(3) equivariant MPNNs were able to achieve an unprecedented accuracy compared to the previous invariant architectures Batzner et al. (2022); Batatia et al. (2022a). It has been recognized that the spherical harmonics expansion and the subsequent Clebsch-Gordan tensor contraction is a bottleneck in the speed and the simplicity of ACE and these equivariant MPNNs. Clever methods have been developed to alleviate this, such as evaluate spherical harmonics in Cartesian space Drautz (2020); Lysogorskiy et al. (2021), absorbing Clebsch Gordan matrix in expansion coefficients during the evaluation stage Lysogorskiy et al. (2021), and using Gaunt coefficients instead of Clebsch-Gordan Luo et al. (2024). There have been other approaches that only operates in the Cartesian space, including NewtonNet Haghighatlari et al. (2022), EGNN Satorras et al. (2021), and PaiNN Schütt et al. (2021) that uses atomic displacement vectors in message passing, the PET based on local coordinate systems Pozdnyakov and Ceriotti (2023), and TensorNet that uses Cartesian tensors to represent pairwise displacements Simeon and De Fabritiis (2023). CACE is different from these approaches in that it can be directly mapped to ACE, so like ACE, it is also complete and body-ordered. The completeness helps the accuracy and the smoothness of the potential, particularly when there exists atomic environments that are degenerate using three-body descriptors or bispectrum Pozdnyakov et al. (2020); Nigam et al. (2023). The body order makes the potential more physically interpretable and may help extrapolation. The angular basis of CACE (Eqn. (2)) is the same as MTP Shapeev (2016), but CACE avoids computing moment tensors, and form a more compact symetric basis set via a simple combination rule that eliminates polynomially depedent features (Fig. 1i). As added advantages, the CACE descriptors are low-dimensional, efficient to evaluate, and are capable of alchemical learning.
@@
A topic that tends to be overlooked in the development of MLIP methods is the stability and generalizability. In practice, a MLIP that is not stable in MD simulations is not useful. Nonetheless, stability and generalization are rarely considered in benchmarks, maybe due to that they are less straightforward to assess than energy and force errors. On the other hand, it has been shown that accuracy of the potential does not imply stability Fu et al. (2023); Stocker et al. (2022). The CACE potential shows high stability and extrapolatability in the datasets we benchmarked: stable MD simulations of water up to 2000 K, MD of ethanol, extrapolation to high temperature and dihedral scan of 3BPA, as well as generalization to the melt and unseen elements in high-entropy alloys. For water, the accuracy of CACE is on par with the most accurate potentials to date, NequIP Batzner et al. (2022) and MACE Batatia et al. (2022a). For small molecules, CACE is less accurate than NequIP or MACE but more accurate than the other recent MLIPs that we compared to. For all practical purposes, the accuracy of CACE for the small molecules is probably enough, considering the error is very small in absolute terms and is way below the chemical accuracy. Moreover, CACE achieved such accuracy with only one message passing layer and relatively few training parameters, and its accuracy can probably be improved with more message passing layers, higher $N_{\mathrm{embedding}}$, $\nu_{\mathrm{max}}$, $l_{\mathrm{max}}$, or a larger MLP at the readout.
@@
We implemented the CACE potential using PyTorch, and the code is available in https://github.com/BingqingCheng/cace. With the current implementation, the training time of the water and the HEA25 datasets took about two days on a vintage GeForce GTX 1080 Ti GPU card, and the small molecules can be trained on a laptop within one or two days. The training should be much faster on state-of-the-art GPUs. For MD simulations of liquid water using the $T=1$ CACE water model on a single Nvidia A100 GPU, systems with 1,536 atoms, 5,184 atoms, 12,288 atoms, 24,000 atoms can run 24 ps per hour, 8 ps per hour, 3.5 ps per hour, 1.7 ps per hour respectively. The ability to run such large system sizes on a single GPU speaks to the low memory consumption of CACE. The implementation of the code can be further optimized: The computation of the node features and the symmetrization step may be compiled and be made faster. The training and the prediction are available on one GPU and can be made more parallelizable. The MD simulations can be performed in the ASE package, but the LAMMPS interface implementation is currently absent. As CACE has a moderate receptive field, so it is in principle possible to make it scale well in MD simulations.
@@
Many engineering aspects of the CACE potential can be further fine-tuned. For example, dataset normalization and internal normalization can have a large effect on training efficiency and outcome Batatia et al. (2022b). Radial basis influences the learning efficiency Bigi et al. (2022), and its best choice is still an open question. Other choices of $F$, $H$ and $G$ functions in the message passing may be used. The training protocol can also influence the outcome.
@@
Other possible future improvements of CACE include pretraining on large datasets before fine-tuning the model. It should be possible to make a foundation model that is generally valid across the periodic table Batatia et al. (2023). The alchemical learning capability of CACE may facilitate this.
V Conclusions
@@
In summary, we propose a new scheme to preserve the rotational symmetry of atomic environments with different body orders, and the operation is entirely in the Cartesian coordinates. The transformation to the spherical harmonic basis, tensor product of atomic basis, and the usage of high-dimensional Clebsch-Gordan coupling matrix are akk avoided. The resulting symmeric features are low-dimensional and polynormially indepedent. As the symmetrization in the Cartesian coordinates is equivalent to the Clebsch-Gordan contraction, the scheme can be applied in other MLIPs that use the ACE framework (e.g. linear ACE Witt et al. (2023), SNAP, SOAP-GAP Bartók et al. (2010), Jacobi-Legendre cluster expansion Domina et al. (2023), ml-ACE Bochkarev et al. (2022), graph ACE Bochkarev et al. (2024)) , or E(3) equivariant message passing neural networks (e.g. NequIP Batzner et al. (2022), MACE Batatia et al. (2022a), BOTNet Batatia et al. (2022b), SEGNN Brandstetter et al. (2021), Allegro Musaelian et al. (2023)).

### v3

@@ I.2 Bulk water
I.3 Small molecules: ethanol and 3BPA
To demonstrate CACE on small organic molecules, we fitted to 1,000 ethanol structures from the MD17 database Chmiela et al. (2017) that is often used for benchmark purposes. Table 2 shows a comparison of the different MLIPs trained on energies and forces for the same MD17-ethanol dataset. The energy accuracy of CACE is similar to the NequIP model with $T=5$, although the CACE force error is larger. The stability (S) in Table 2 measures how long MD simulations at 500 K using the MLIP is physically sensical (no pathological behavior or enter physically prohibitive states) in runs with a maximum length of 300 ps, so 300 ps is the maximum score in this case Fu et al. (2023). The stability values for other MLIPs are from Ref. Fu et al. (2023) which used 10,000 training structures and a time step of 0.5 fs in MD. For CACE, we used 1,000 training structures and a time step of 1 fs, which is more stringent, but the MD had nevertheless remained stable. As observed in Refs. Fu et al. (2023); Stocker et al. (2022), force accuracy does not imply stability, and the latter is a key metric for MLIPs.
@@ I.3 Small molecules: ethanol and 3BPA
Table 2: Mean absolute error (MAE) of energy (E) and force (F) for ethanol in the MD-17 dataset, in units of meV and meVÅ-1, respectively, trained on 1000 reference configurations. The stability (S) in the unit of ps shows how long the simulation using the MLIP is physically realistic in MD of a maximum length of 300 ps at 500 K. Besides CACE, the errors are obtained from respective references, and the stability is from Ref. Fu et al. (2023).
@@
Table 3: Root mean square errors (RMSE) for energy (E) and force (F) on the 3BPA temperature transferability dataset, reported in units of meV and meVÅ-1. All models were trained on $T=300$ K data. Results of previous models except for NequIP and MACE are from Kovács et al. (2021). ANI-2x was trained on 8.9 million structures, and ANI-pretrained was pretrained and fine-tuned on this dataset.
Figure 4: The dihedral scan benchmark of the 3BPA molecule. a Molecular structure and three dihedral angles. b The dihedral potential energy landscape for $\beta=120^{\circ}$, as predicted by DFT and CACE. The white dots on the DFT potential energy surface correspond to all the training configurations that have $\beta$ between 100 and 140.
As another example, we fitted to the 3BPA (3-(benzyloxy)pyridin-2-amine) dataset Kovács et al. (2021). Table 3 shows the errors of different potentials when extrapolating to higher temperature. NequIP and MACE perform the best, while CACE is somewhere between them and ACE. We then compared the dihedral potential energy surface of the molecule. Fig. 4 shows the complex energy landscapes of the $\alpha$-$\gamma$ plane predicted by DFT and by CACE, for the case of $\beta=120^{\circ}$, the plane with the fewest training points. Nevertheless, the CACE landscape closely resembles the ground truth and is regular, more so than the other potentials benchmarked in Ref. Kovács et al. (2021) despite those were trained on a more diverse dataset including high temperature configurations.
I.4 25-element high-entropy alloys
High-entropy alloys, which are composed of several metallic elements, have unique mechanical, thermodynamics, and catalytic properties George et al. (2019). They are challenging to model due to the high number of elements involved, for which many MLIPs scale poorly with. Ref. Lopanitsyna et al. (2023) introduced a 25 d-block transition metal HEA dataset of distorted crystalline structures. Ref. Lopanitsyna et al. (2023) further performed an alchemical learning (AL) that explicit compresses chemical space, leading to a model that is both interpretable and very stable.
@@
Table 4: Mean absolute error (MAE) of energy (E) and force (F) for the HEA25 dataset, reported in units of meV and meVÅ-1. There are three sets of benchmarks on test error, alchemical extrapolation to unseen Re and Os substitute atoms, and temperature transferability to 5000 K.
Figure 5: Benchmark on the HEA25 dataset. a Learning curves for different models. Both alchemical learning (AL) models are from Ref. Lopanitsyna et al. (2023). b The first two principal components (PCs) of the CACE embedding matrix. The periods are highlighted with orange, blue, and green lines. Interpolated positions for Re and Os are indicated with empty circles. c Energy comparison between DFT and CACE for a high temperature (5000 K) and a low temperature (300 K) trajectory. The inset shows example solid and melt configurations from the 300 K and the 5000 K trajectories, respectively. d Comparison between the substitution energy $E_{\mathrm{sub}}$, the potential energy difference between the original structure and the structure substituted with Re and Os element per substitution atom, computed using DFT and CACE. The inset shows the parity plot for the force components computed for those structures.
@@
A stringent way to demonstrate alchemical learning is by extrapolating to unseen elements. In this case, Re and Os are absent from the training set. Ref. Lopanitsyna et al. (2023) provides a hold-out set containing 61 pairs of structures, each of 36 or 48 atoms. Each pair has one fcc or bcc structure with only the original 25 elements, and another structure with up to 11 atoms randomly substituted with Re and Os. We simply took the CACE model trained for the 25-element dataset and obtained the chemical embedding of Re and Os by a linear interpolation between W and Ir, as illustrated by the empty circles in Fig. 5b. We also added atomic-energy baselines for Re and Os (obtained by fitting the residual energy to a two-parameter linear model that depends only on the Re and Os content). In Fig. 5d we show the parity plot for the substitution energy, defined as the potential energy difference between the original and the substituted structure per substitution atom. Impressively, the CACE error for those substituted structures with unseen elements is rather similar to the test error on the 25-element training set, and is lower than the AL method in Ref. Lopanitsyna et al. (2023). The force comparison, shown in the inset of Fig. 5d, shows the same trend. With the PET potential, it is not clear whether alchemical learning can be performed in a straightforward way.
@@ II Discussion
There have been other MLIP approaches that only operate in Cartesian space, including NewtonNet Haghighatlari et al. (2022), EGNN Satorras et al. (2021), and PaiNN Schütt et al. (2021) that use atomic displacement vectors in message passing, the PET based on local coordinate systems Pozdnyakov and Ceriotti (2023), and TensorNet that uses Cartesian tensors to represent pairwise displacements Simeon and De Fabritiis (2023). CACE is different from these approaches in that it can be directly mapped to ACE, so like ACE, it is also complete and body-ordered. The completeness helps the accuracy and the smoothness of the potential, particularly when there exist atomic environments that are degenerate using three-body descriptors or bispectrum Pozdnyakov et al. (2020); Nigam et al. (2023). The body order makes the potential more physically interpretable and may help extrapolation. The angular basis of CACE (Eqn. (2)) is the same as MTP Shapeev (2016), but CACE avoids computing moment tensors, and form a more compact symetric basis set via a simple combination rule that eliminates polynomially depedent features (Fig. 1i). As added advantages, the CACE descriptors have optimized radial channel coupling, and are capable of alchemical learning.
@@
A topic that tends to be overlooked in the development of MLIP methods is the stability and generalizability. In practice, a MLIP that is not stable in MD simulations is not useful. Nonetheless, stability and generalization are rarely considered in benchmarks, maybe due to that they are less straightforward to assess than energy and force errors. On the other hand, it has been shown that accuracy of the potential does not imply stability Fu et al. (2023); Stocker et al. (2022). The CACE potential shows high stability and extrapolatability in the datasets we benchmarked: stable MD simulations of water up to 2000 K, MD of ethanol, extrapolation to high temperature and dihedral scan of 3BPA, as well as generalization to the melt and unseen elements in high-entropy alloys. For water, the accuracy of CACE is on par with the most accurate potentials to date, NequIP Batzner et al. (2022) and MACE Batatia et al. (2022a). For small molecules, CACE is less accurate than NequIP or MACE but more accurate than the other recent MLIPs that we compared to. For all practical purposes, the accuracy of CACE for the small molecules is probably enough, considering that the error is very small in absolute terms and is much lower than the chemical accuracy. Moreover, CACE achieved such accuracy with only one message passing layer and relatively few training parameters, and its accuracy can probably be improved with more message passing layers, higher $N_{\mathrm{embedding}}$, $\nu_{\mathrm{max}}$, $l_{\mathrm{max}}$, or a larger MLP at the readout.
NequiIP Batzner et al. (2022) MACE Batatia et al. (2022a) CACE
$4.5$, $T=3$ $6$ $5.5$, $T=1$
1,536 atoms 20 ps per h 18 ps per h 24 ps per h
5,184 atoms 7 ps per h 4.8 ps per h 8 ps per h
12,288 atoms 3 ps per h 2 ps per h 3.5 ps per h
24,000 atoms 1.65 ps per h N.A. 1.7 ps per h
Table 5: Speed of MD of liquid water (in picosecond of simulation per hour) using different water models on a single Nvidia A100 GPU.
@@
We implemented the CACE potential using PyTorch, and the code is available in https://github.com/BingqingCheng/cace. With the current implementation, the training time of the water and the HEA25 datasets took about two days on a vintage GeForce GTX 1080 Ti GPU card, and the small molecules can be trained on a laptop within one or two days. The training should be much faster on state-of-the-art GPUs. For the speed of MD, Table 5 compares the timing on one Nvidia A100 GPU using three water models in Table 1. Note that such comparison depends very much on the specific models used (higher $\nu$, $l_{\mathrm{max}}$, $r_{\mathrm{cut}}$, and $T$ all add to the cost) as well as the implementation. CACE is marginally faster in this comparison and can drive simulations of 24,000 atoms. The ability to run such large system sizes on a single GPU speaks to the low memory consumption of CACE. The implementation of the code can be further optimized: The computation of the node features and the symmetrization step may be compiled and be made faster. The training and the prediction are available on one GPU and can be made more parallelizable. The MD simulations can be performed in the ASE package, but the LAMMPS interface implementation is currently absent. As CACE has a moderate receptive field, so it is in principle possible to make it scale well in MD simulations.
@@
Many engineering aspects of the CACE potential can be further fine-tuned. For example, dataset normalization and internal normalization can have a large effect on training efficiency and outcome Batatia et al. (2022b). Radial basis influences the learning efficiency Bigi et al. (2022), and its best choice is still an open question. Other choices of the $F$, $H$ and $G$ functions in the message passing may be used. One can also use the product of $B_{i}^{(t)})$ and $\chi_{cn\mathbf{l}}$ in the message function, exploiting the relationship in Eqn. (3). The training protocol can also influence the outcome.
@@
Other possible future improvements of CACE include pretraining on large datasets before fine-tuning the model. It should be possible to develop a foundation model that is generally valid in the periodic table Batatia et al. (2023). The alchemical learning capacity of CACE may facilitate this.
@@
In summary, we propose a scheme to preserve the rotational symmetry of atomic environments with different body orders, and the operation is entirely in the Cartesian coordinates. The transformation to the spherical harmonic basis, tensor product of atomic basis, and the usage of high-dimensional Clebsch-Gordan coupling matrix are all avoided. The resulting symmeric features are low-dimensional and polynormially indepedent. As the symmetrization in the Cartesian coordinates is equivalent to the Clebsch-Gordan contraction, the scheme can be applied in other MLIPs that use the ACE framework (e.g. linear ACE Witt et al. (2023), SNAP, SOAP-GAP Bartók et al. (2010), Jacobi-Legendre cluster expansion Domina et al. (2023), ml-ACE Bochkarev et al. (2022), graph ACE Bochkarev et al. (2024)) , or E(3) equivariant message passing neural networks (e.g. NequIP Batzner et al. (2022), MACE Batatia et al. (2022a), BOTNet Batatia et al. (2022b), SEGNN Brandstetter et al. (2021), Allegro Musaelian et al. (2023)).
III Methods
Bulk water The water dataset has 1,593 liquid water configurations, each of 64 molecules Cheng et al. (2019). It was computed at the revPBE0-D3 level of density functional theory (DFT). 90% of the data are used in the training. We used a cutoff $r_{\mathrm{cut}}=5.5$, 6 Bessel radial functions and $c=12$, $l_{\mathrm{max}}=3$, $\nu_{\mathrm{max}}=3$, $N_{\mathrm{embedding}}=3$, and no message passing ($T=0$) or one message passing layer ($T=1$). The fitted CACE models have relatively few parameters: 24,572 trainable parameters for the CACE water model with $T=0$, and 69,320 trainable parameters for the model with $T=1$.
In Table 1, the NequIP and DeepMD models are trained by us, and the others are from the references. In addition, we also trained a MACE model with E RMSE=1.9 meV/H2O, F RMSE=46.8 meVÅ-1. as the original Behler-Parrinello neural network (BPNN) in Ref. Cheng et al. (2019) was trained using the RuNNer code Behler (2018), we also trained another one using N2P2 Singraber et al. (2019) and obtained E RMSE=5.8 meV/H2O, F RMSE=108 meVÅ-1.
Small molecules The ethanol data from the MD17 database Chmiela et al. (2017) were sampled from DFT molecular dynamics simulations at 500 K. We used a cutoff $r_{\mathrm{cut}}=4$, 6 Bessel radial functions, $c=12$, $l_{\mathrm{max}}=4$, $\nu_{\mathrm{max}}=3$, $N_{\mathrm{embedding}}=4$, and one message passing layer ($T=1$). The model was trained on 1,000 structures randomly selected from the dataset (90% train/10% validation split), and tested on another 10,000 random structures.
The 3BPA dataset Kovács et al. (2021) set consists of 500 training structures at 300 K. Test data are from simulations at 300 K, 600 K, and 1200 K. For fitting the CACE model, we used $r_{\mathrm{cut}}=6$, 8 Bessel radial functions, $c=12$, $l_{\mathrm{max}}=4$, $\nu_{\mathrm{max}}=3$, $N_{\mathrm{embedding}}=4$, and $T=1$.
25-element high-entropy alloys The HEA dataset from Ref. Lopanitsyna et al. (2023) consists of 25,630 distorted crystalline structures containing 36 or 48 atoms on bcc or fcc lattices. We took a test set containing 1,000 configurations, and used up to the rest of the 24,630 data points for training and validation. We used a CACE model with $r_{\mathrm{cut}}=4.5$, $l_{\mathrm{max}}=3$, $\nu_{\mathrm{max}}=3$, $N_{\mathrm{embedding}}=3$, and $T=1$.